Let a > 1. Suppose S = {2, 3,...,} and P(k) = a4k-1 , k = 2, 3...

Find P(outcome is less than or equal to 4). Express the answer in terms of a.
Answer #1

for this to be valid: \sumP(k) should be equal to 1

\sumP(k) =c(1/a7 +1/a11+1/a15+a19+... ) =c*(1/a7)/(1-1/a4) =(c/a3)*(1/(a4-1)) =1

c=(a4-1)*a3
hence  P(outcome is less than or equal to 4)=P(X=2)+P(X=3)+P(X=4)

=a3*(a4-1)*(1/a7 +1/a11+1/a15) =a3*(a4-1)*(a8+a4+1)/a15 =(a4-1)*(a8+a4+1)/a12